# New results for the Cahn-Hilliard equation with non-degenerate mobility: well-posedness and longtime behavior

### v1

@@ 2.1. Notation and function spaces
$\nabla\mathcal{G}_{\varphi_{1}^{\varepsilon}}\partial_{t}\Phi\to\nabla\mathcal{G}_{\varphi_{1}}\partial_{t}\Phi\ \text{strongly in }L^{2}(0,T;L^{2}(\Omega)),$ (3.34)
@@
$\nabla\mathcal{G}_{\varphi_{1}^{\varepsilon}}\Phi\to\nabla\mathcal{G}_{\varphi_{1}}\Phi\ \text{strongly in }L^{2}(0,T;L^{2}(\Omega)),$ (3.35)
@@
$\leq C\|\nabla\varphi_{1}^{\varepsilon}\|_{L^{6}(\Omega)}\|\nabla\mathcal{G}_{\varphi_{1}^{\varepsilon}}\Phi\|_{L^{6}(\Omega)}^{2}+C\|D^{2}\mathcal{G}_{\varphi_{1}^{\varepsilon}}\Phi\|_{L^{4}(\Omega)}\|\nabla\mathcal{G}_{\varphi_{1}^{\varepsilon}}\Phi\|_{L^{4}(\Omega)}$
$\leq C\|\nabla\varphi_{1}^{\varepsilon}\|_{L^{2}(\Omega)}^{\frac{1}{3}}\|\nabla\varphi_{1}^{\varepsilon}\|_{H^{1}(\Omega)}^{\frac{2}{3}}\|\nabla\mathcal{G}_{\varphi_{1}^{\varepsilon}}\Phi\|_{L^{2}(\Omega)}^{\frac{2}{3}}\|\nabla\mathcal{G}_{\varphi_{1}^{\varepsilon}}\Phi\|_{H^{1}(\Omega)}^{\frac{4}{3}}$
$\quad+C\|D^{2}\mathcal{G}_{\varphi_{1}^{\varepsilon}}\Phi\|_{L^{4}(\Omega)}\|\nabla\mathcal{G}_{\varphi_{1}^{\varepsilon}}\Phi\|_{L^{2}(\Omega)}^{\frac{1}{2}}\|\nabla\mathcal{G}_{\varphi_{1}^{\varepsilon}}\Phi\|_{H^{1}(\Omega)}^{\frac{1}{2}}.$
@@
$\|\nabla\left(b^{\prime}(\varphi_{1}^{\varepsilon})|\nabla\mathcal{G}_{\varphi_{1}^{\varepsilon}}\Phi|^{2}\right)\|_{L^{2}(\Omega)}$ $\leq C\|\nabla\varphi_{1}^{\varepsilon}\|_{H^{1}(\Omega)}^{\frac{2}{3}}\|\nabla\mathcal{G}_{\varphi_{1}^{\varepsilon}}\Phi\|_{H^{1}(\Omega)}^{\frac{4}{3}}+C\|D^{2}\mathcal{G}_{\varphi_{1}^{\varepsilon}}\Phi\|_{L^{4}(\Omega)}\|\nabla\mathcal{G}_{\varphi_{1}^{\varepsilon}}\Phi\|_{H^{1}(\Omega)}^{\frac{1}{2}}$
$\leq C\|\nabla\varphi_{1}^{\varepsilon}\|_{H^{1}(\Omega)}^{2}+C\|\nabla\mathcal{G}_{\varphi_{1}^{\varepsilon}}\Phi\|_{H^{1}(\Omega)}^{2}+C\|D^{2}\mathcal{G}_{\varphi_{1}^{\varepsilon}}\Phi\|_{L^{4}(\Omega)}^{\frac{4}{3}},$
@@
which entails that $g_{1}=b^{\prime\prime}(\varphi_{1})\nabla\varphi_{1}|\nabla\mathcal{G}_{\varphi_{1}}\Phi|^{2}$. In a similar matter, owing to (3.35), (3.37), and observing that $b^{\prime}(\varphi_{1}^{\varepsilon})\to b^{\prime}(\varphi_{1})$ strongly in $L^{8}(0,T;L^{4}(\Omega))$, we obtain
@@
This allows us to conclude that $g_{2}=2b^{\prime}(\varphi_{1})D^{2}\mathcal{G}_{\varphi_{1}}\Phi\nabla\mathcal{G}_{\varphi_{1}}\Phi$. Therefore, by exploiting (3.43) and (3.44) in (3.42), we deduce that
@@
$\|\mathcal{G}_{\varphi_{1}}\Phi\|_{H^{2}(\Omega)}\leq C\left(\|\nabla\varphi_{1}\|_{L^{2}(\Omega)}\|\varphi_{1}\|_{H^{2}(\Omega)}\|\nabla\mathcal{G}_{\varphi_{1}}\Phi\|_{L^{2}(\Omega)}+\|\nabla\mathcal{G}_{\varphi_{1}}\Phi\|_{L^{2}(\Omega)}^{\frac{1}{2}}\|\nabla\Phi\|_{L^{2}(\Omega)}^{\frac{1}{2}}\right)\!.$ (3.46)
@@
$\leq C\left(\|\nabla\varphi_{1}\|_{L^{2}(\Omega)}^{\frac{1}{2}}\|\varphi_{1}\|_{H^{2}(\Omega)}^{\frac{1}{2}}\|\nabla\mathcal{G}_{\varphi_{1}}\Phi\|_{L^{2}(\Omega)}+\|\nabla\mathcal{G}_{\varphi_{1}}\Phi\|_{L^{2}(\Omega)}^{\frac{3}{4}}\|\nabla\Phi\|_{L^{2}(\Omega)}^{\frac{1}{4}}\right)\!.$
@@
$\left\|\nabla\mathcal{G}_{\varphi_{1}}\Phi\right\|_{L^{4}(\Omega)}\leq C\left(\|\varphi_{1}\|_{H^{2}(\Omega)}^{\frac{1}{2}}\|\nabla\mathcal{G}_{\varphi_{1}}\Phi\|_{L^{2}(\Omega)}+\|\nabla\mathcal{G}_{\varphi_{1}}\Phi\|_{L^{2}(\Omega)}^{\frac{3}{4}}\|\nabla\Phi\|_{L^{2}(\Omega)}^{\frac{1}{4}}\right)\!.$ (3.48)
@@
$\|\mathcal{G}_{\varphi_{1}}\Phi\|_{W^{2,4}(\Omega)}$ $\leq C\left(\left\|\nabla\varphi_{1}\right\|_{L^{2}(\Omega)}^{\frac{1}{4}}\left\|\varphi_{1}\right\|_{H^{2}(\Omega)}^{\frac{3}{4}}\left\|\nabla\mathcal{G}_{\varphi_{1}}\Phi\right\|_{L^{2}(\Omega)}^{\frac{1}{4}}\left\|\mathcal{G}_{\varphi_{1}}\Phi\right\|_{H^{2}(\Omega)}^{\frac{3}{4}}+\|\Phi\|_{L^{2}(\Omega)}^{\frac{1}{2}}\|\nabla\Phi\|_{L^{2}(\Omega)}^{\frac{1}{2}}\right)$
@@
$\|\mathcal{G}_{\varphi_{1}}\Phi\|_{W^{2,4}(\Omega)}$ $\leq C\left\|\varphi_{1}\right\|_{H^{2}(\Omega)}^{\frac{3}{2}}\left\|\nabla\mathcal{G}_{\varphi_{1}}\Phi\right\|_{L^{2}(\Omega)}+C\left\|\varphi_{1}\right\|_{H^{2}(\Omega)}^{\frac{3}{4}}\left\|\nabla\mathcal{G}_{\varphi_{1}}\Phi\right\|_{L^{2}(\Omega)}^{\frac{5}{8}}\left\|\nabla\Phi\right\|_{L^{2}(\Omega)}^{\frac{3}{8}}$
$\quad+C\|\nabla\mathcal{G}_{\varphi_{1}}\Phi\|_{L^{2}(\Omega)}^{\frac{1}{4}}\|\nabla\Phi\|_{L^{2}(\Omega)}^{\frac{3}{4}}.$ (3.49)
@@
Concerning $I_{3}$, we observe from (2.11) and (1.18a) that
@@
$=\int_{\Omega}\nabla\mu_{2}\cdot\left(b(\varphi_{2})-b(\varphi_{1})\right)\nabla\mathcal{G}_{\varphi_{1}}\Phi\,{\rm d}x+\int_{\Omega}\nabla\mu_{2}\cdot b(\varphi_{1})\nabla\mathcal{G}_{\varphi_{1}}\Phi\,{\rm d}x-\int_{\Omega}\nabla\mu_{2}\cdot b(\varphi_{2})\nabla\mathcal{G}_{\varphi_{2}}\Phi\,{\rm d}x$
@@
$\leq C\left\|\nabla\mu_{2}\right\|_{L^{2}(\Omega)}\|\Phi\|_{L^{4}(\Omega)}\left\|\nabla\mathcal{G}_{\varphi_{1}}\Phi\right\|_{L^{4}(\Omega)}$
$\leq C\left\|\nabla\mu_{2}\right\|_{L^{2}(\Omega)}\|\nabla\mathcal{G}_{\varphi_{1}}\Phi\|_{L^{2}(\Omega)}^{\frac{1}{4}}\|\nabla\Phi\|_{L^{2}(\Omega)}^{\frac{3}{4}}$
×$\quad\times\left(\|\varphi_{1}\|_{H^{2}(\Omega)}^{\frac{1}{2}}\|\nabla\mathcal{G}_{\varphi_{1}}\Phi\|_{L^{2}(\Omega)}+\|\nabla\mathcal{G}_{\varphi_{1}}\Phi\|_{L^{2}(\Omega)}^{\frac{3}{4}}\|\nabla\Phi\|_{L^{2}(\Omega)}^{\frac{1}{4}}\right)$
$\leq C\left\|\nabla\mu_{2}\right\|_{L^{2}(\Omega)}\|\varphi_{1}\|_{H^{2}(\Omega)}^{\frac{1}{2}}\|\nabla\mathcal{G}_{\varphi_{1}}\Phi\|_{L^{2}(\Omega)}^{\frac{5}{4}}\|\nabla\Phi\|_{L^{2}(\Omega)}^{\frac{3}{4}}$
$\quad+C\left\|\nabla\mu_{2}\right\|_{L^{2}(\Omega)}\|\nabla\mathcal{G}_{\varphi_{1}}\Phi\|_{L^{2}(\Omega)}\|\nabla\Phi\|_{L^{2}(\Omega)}$
$\leq\frac{1}{8}\|\nabla\Phi\|_{L^{2}(\Omega)}^{2}+C\left(\left\|\nabla\mu_{2}\right\|_{L^{2}(\Omega)}^{\frac{8}{5}}\|\varphi_{1}\|_{H^{2}(\Omega)}^{\frac{4}{5}}+\left\|\nabla\mu_{2}\right\|_{L^{2}(\Omega)}^{2}\right)\|\nabla\mathcal{G}_{\varphi_{1}}\Phi\|_{L^{2}(\Omega)}^{2}$
$\leq\frac{1}{8}\|\nabla\Phi\|_{L^{2}(\Omega)}^{2}+C\left(\left\|\nabla\mu_{2}\right\|_{L^{2}(\Omega)}^{2}+\|\varphi_{1}\|_{H^{2}(\Omega)}^{4}\right)\|\nabla\mathcal{G}_{\varphi_{1}}\Phi\|_{L^{2}(\Omega)}^{2}.$ (3.51)
@@
$\frac{{\rm d}}{{\rm d}t}\left\|\sqrt{b(\varphi_{1})}\nabla\mathcal{G}_{\varphi_{1}}\Phi\right\|_{L^{2}(\Omega)}^{2}+\left\|\nabla\Phi\right\|_{L^{2}(\Omega)}^{2}\leq h(t)\left\|\sqrt{b(\varphi_{1})}\nabla\mathcal{G}_{\varphi_{1}}\Phi\right\|_{L^{2}(\Omega)}^{2}\!.$ (3.53)
@@
$\begin{split}&\varphi\in L^{\infty}(0,\infty;W^{2,p}(\Omega)),\quad\partial_{t}\varphi\in L_{\operatorname{uloc}}^{2}(0,\infty;H^{1}(\Omega)),\\ &\varphi\in L^{\infty}(\Omega\times(0,\infty))\quad\text{such that}\quad|\varphi(x,t)|<1\ \text{a.e. }(x,t)\in\Omega\times(0,\infty),\\ &\mu\in L^{\infty}(0,\infty;H^{1}(\Omega))\cap L_{\operatorname{uloc}}^{4}([0,\infty);H^{2}(\Omega)),\quad F^{\prime}(\phi)\in L^{\infty}(0,\infty;L^{p}(\Omega)),\end{split}$ (4.1)
@@
First of all, owing to [5, Theorem 2.2] (cf. the three dimensional case), there exists a unique strong solution $\varphi$ to problem (1.1a)-(1.1b) on a maximal time interval $[0,T_{\star})$. In particular, the solution $\varphi$ is the limit of an approximating sequence $\{\varphi_{\varepsilon}\}$, where $\varphi_{\varepsilon}$ is the solution to a suitable approximation problem associated with a regularized potential $F_{\varepsilon}$. Our aim is showing that $T_{\star}=+\infty$. More precisely, we will show that $\varphi$ satisfies (4.1). To this end, we will perform global-in-time estimates leading to (4.2)-(4.4) by directly working on the solution $\varphi$. However, a rigorous proof should be carried out by making the computations below at the approximation level $\varphi_{\varepsilon}$, where also the initial datum is properly regularized (see the proof of [28, Theorem 4.1]), and by passing ultimately to the limit. Since this procedure is nowadays well established in the literature, we will leave this step to the interested reader.
@@
almost everywhere on $(0,T_{\star})$. The way to handle the first term on the right hand side in (4.8) is well known (see (4.14) below). Instead, we need to find a suitable control for the second term in the right-hand side. In order to do so we present two different approaches.
@@
$\left\|\mathcal{G}_{\varphi}\partial_{t}\varphi\right\|_{H^{2}(\Omega)}\leq C^{\frac{1}{1-\tau}}\left(\frac{2}{\tau(1-\tau)}\right)^{\frac{1}{2(1-\tau)}}\left\|\nabla\varphi\right\|_{L^{2}(\Omega)}^{\frac{\tau}{1-\tau}}\left\|\varphi\right\|_{H^{2}(\Omega)}\left\|\nabla\mathcal{G}_{\varphi}\partial_{t}\varphi\right\|_{L^{2}(\Omega)}+\frac{C}{1-\tau}\left\|\partial_{t}\varphi\right\|_{L^{2}(\Omega)}\!.$
@@
$\begin{split}\left\|\nabla\mu\right\|^{4}_{L^{2(1+\tau)}(\Omega)}&\leq C(1+\tau)^{2}\|\nabla\mu\|_{L^{2}(\Omega)}^{\frac{4}{1+\tau}}\|\mathcal{G}_{\varphi}\partial_{t}\varphi\|_{H^{2}(\Omega)}^{\frac{4\tau}{1+\tau}}\\ &\leq C(1+\tau)^{2}\left(\frac{1}{\tau}\right)^{\frac{5\tau}{2(1+\tau)}}\|\nabla\varphi\|_{L^{2}(\Omega)}^{\frac{4\tau^{2}}{1-\tau^{2}}}\|\varphi\|_{H^{2}(\Omega)}^{\frac{4\tau}{1+\tau}}\|\nabla\mu\|_{L^{2}(\Omega)}^{4}\\ &\quad+C(1+\tau)^{2}\|\nabla\mu\|_{L^{2}(\Omega)}^{\frac{4}{1+\tau}}\|\partial_{t}\varphi\|_{L^{2}(\Omega)}^{\frac{4\tau}{1+\tau}}\\ &\leq C(1+\tau)^{2}\left(\frac{1}{\tau}\right)^{\frac{5\tau}{2(1+\tau)}}\|\nabla\varphi\|_{L^{2}(\Omega)}^{\frac{4\tau^{2}}{1-\tau^{2}}}\|\varphi\|_{H^{2}(\Omega)}^{\frac{4\tau}{1+\tau}}\|\nabla\mu\|_{L^{2}(\Omega)}^{4}\\ &\quad+\frac{2\tau}{1+\tau}\|\partial_{t}\varphi\|_{L^{2}(\Omega)}^{2}+\left(\frac{1-\tau}{1+\tau}\right)\left[C(1+\tau)^{2}\right]^{\frac{1+\tau}{1-\tau}}\ \|\nabla\mu\|_{L^{2}(\Omega)}^{\frac{4}{1-\tau}}.\end{split}$

### v2

@@ 2.1. Notation and function spaces
$\nabla\mathcal{G}_{\varphi_{1}^{\varepsilon}}\partial_{t}\Phi\to\nabla\mathcal{G}_{\varphi_{1}}\partial_{t}\Phi\quad\text{strongly in }L^{2}(0,T;L^{2}(\Omega)),$ (3.34)
@@
$\nabla\mathcal{G}_{\varphi_{1}^{\varepsilon}}\Phi\to\nabla\mathcal{G}_{\varphi_{1}}\Phi\quad\text{strongly in }L^{2}(0,T;L^{2}(\Omega)),$ (3.35)
@@
$\leq C\left\|\nabla\varphi_{1}^{\varepsilon}\right\|_{L^{6}(\Omega)}\left\|\nabla\mathcal{G}_{\varphi_{1}^{\varepsilon}}\Phi\right\|_{L^{6}(\Omega)}^{2}+C\left\|D^{2}\mathcal{G}_{\varphi_{1}^{\varepsilon}}\Phi\right\|_{L^{4}(\Omega)}\left\|\nabla\mathcal{G}_{\varphi_{1}^{\varepsilon}}\Phi\right\|_{L^{4}(\Omega)}$
$\leq C\left\|\nabla\varphi_{1}^{\varepsilon}\right\|_{L^{2}(\Omega)}^{\frac{1}{3}}\left\|\nabla\varphi_{1}^{\varepsilon}\right\|_{H^{1}(\Omega)}^{\frac{2}{3}}\left\|\nabla\mathcal{G}_{\varphi_{1}^{\varepsilon}}\Phi\right\|_{L^{2}(\Omega)}^{\frac{2}{3}}\left\|\nabla\mathcal{G}_{\varphi_{1}^{\varepsilon}}\Phi\right\|_{H^{1}(\Omega)}^{\frac{4}{3}}$
$\quad+C\left\|D^{2}\mathcal{G}_{\varphi_{1}^{\varepsilon}}\Phi\right\|_{L^{4}(\Omega)}\left\|\nabla\mathcal{G}_{\varphi_{1}^{\varepsilon}}\Phi\right\|_{L^{2}(\Omega)}^{\frac{1}{2}}\left\|\nabla\mathcal{G}_{\varphi_{1}^{\varepsilon}}\Phi\right\|_{H^{1}(\Omega)}^{\frac{1}{2}}.$
@@
$\left\|\nabla\left(b^{\prime}(\varphi_{1}^{\varepsilon})|\nabla\mathcal{G}_{\varphi_{1}^{\varepsilon}}\Phi|^{2}\right)\right\|_{L^{2}(\Omega)}$ $\leq C\left\|\nabla\varphi_{1}^{\varepsilon}\right\|_{H^{1}(\Omega)}^{\frac{2}{3}}\left\|\nabla\mathcal{G}_{\varphi_{1}^{\varepsilon}}\Phi\right\|_{H^{1}(\Omega)}^{\frac{4}{3}}+C\left\|D^{2}\mathcal{G}_{\varphi_{1}^{\varepsilon}}\Phi\right\|_{L^{4}(\Omega)}\left\|\nabla\mathcal{G}_{\varphi_{1}^{\varepsilon}}\Phi\right\|_{H^{1}(\Omega)}^{\frac{1}{2}}$
$\leq C\left\|\nabla\varphi_{1}^{\varepsilon}\right\|_{H^{1}(\Omega)}^{2}+C\left\|\nabla\mathcal{G}_{\varphi_{1}^{\varepsilon}}\Phi\right\|_{H^{1}(\Omega)}^{2}+C\left\|D^{2}\mathcal{G}_{\varphi_{1}^{\varepsilon}}\Phi\right\|_{L^{4}(\Omega)}^{\frac{4}{3}},$
@@
which entails that
$g_{1}=b^{\prime\prime}(\varphi_{1})\nabla\varphi_{1}|\nabla\mathcal{G}_{\varphi_{1}}\Phi|^{2}.$
In a similar matter, owing to (3.35), (3.37), and observing that $b^{\prime}(\varphi_{1}^{\varepsilon})\to b^{\prime}(\varphi_{1})$ strongly in $L^{8}(0,T;L^{4}(\Omega))$, we obtain
@@
This allows us to conclude that
$g_{2}=2b^{\prime}(\varphi_{1})D^{2}\mathcal{G}_{\varphi_{1}}\Phi\nabla\mathcal{G}_{\varphi_{1}}\Phi.$
Therefore, by exploiting (3.43) and (3.44) in (3.42), we deduce that
@@
$\left\|\mathcal{G}_{\varphi_{1}}\Phi\right\|_{H^{2}(\Omega)}\leq C\left(\left\|\nabla\varphi_{1}\right\|_{L^{2}(\Omega)}\left\|\varphi_{1}\right\|_{H^{2}(\Omega)}\left\|\nabla\mathcal{G}_{\varphi_{1}}\Phi\right\|_{L^{2}(\Omega)}+\left\|\nabla\mathcal{G}_{\varphi_{1}}\Phi\right\|_{L^{2}(\Omega)}^{\frac{1}{2}}\left\|\nabla\Phi\right\|_{L^{2}(\Omega)}^{\frac{1}{2}}\right)\!.$ (3.46)
@@
$\leq C\left(\|\nabla\varphi_{1}\|_{L^{2}(\Omega)}^{\frac{1}{2}}\|\varphi_{1}\|_{H^{2}(\Omega)}^{\frac{1}{2}}\left\|\nabla\mathcal{G}_{\varphi_{1}}\Phi\right\|_{L^{2}(\Omega)}+\left\|\nabla\mathcal{G}_{\varphi_{1}}\Phi\right\|_{L^{2}(\Omega)}^{\frac{3}{4}}\left\|\nabla\Phi\right\|_{L^{2}(\Omega)}^{\frac{1}{4}}\right)\!.$
@@
$\left\|\nabla\mathcal{G}_{\varphi_{1}}\Phi\right\|_{L^{4}(\Omega)}\leq C\left(\|\varphi_{1}\|_{H^{2}(\Omega)}^{\frac{1}{2}}\left\|\nabla\mathcal{G}_{\varphi_{1}}\Phi\right\|_{L^{2}(\Omega)}+\left\|\nabla\mathcal{G}_{\varphi_{1}}\Phi\right\|_{L^{2}(\Omega)}^{\frac{3}{4}}\left\|\nabla\Phi\right\|_{L^{2}(\Omega)}^{\frac{1}{4}}\right)\!.$ (3.48)
@@
$\left\|\mathcal{G}_{\varphi_{1}}\Phi\right\|_{W^{2,4}(\Omega)}$ $\leq C\left(\left\|\nabla\varphi_{1}\right\|_{L^{2}(\Omega)}^{\frac{1}{4}}\left\|\varphi_{1}\right\|_{H^{2}(\Omega)}^{\frac{3}{4}}\left\|\nabla\mathcal{G}_{\varphi_{1}}\Phi\right\|_{L^{2}(\Omega)}^{\frac{1}{4}}\left\|\mathcal{G}_{\varphi_{1}}\Phi\right\|_{H^{2}(\Omega)}^{\frac{3}{4}}+\left\|\Phi\right\|_{L^{2}(\Omega)}^{\frac{1}{2}}\left\|\nabla\Phi\right\|_{L^{2}(\Omega)}^{\frac{1}{2}}\right)$
@@
$\left\|\mathcal{G}_{\varphi_{1}}\Phi\right\|_{W^{2,4}(\Omega)}$ $\leq C\left\|\varphi_{1}\right\|_{H^{2}(\Omega)}^{\frac{3}{2}}\left\|\nabla\mathcal{G}_{\varphi_{1}}\Phi\right\|_{L^{2}(\Omega)}+C\left\|\varphi_{1}\right\|_{H^{2}(\Omega)}^{\frac{3}{4}}\left\|\nabla\mathcal{G}_{\varphi_{1}}\Phi\right\|_{L^{2}(\Omega)}^{\frac{5}{8}}\left\|\nabla\Phi\right\|_{L^{2}(\Omega)}^{\frac{3}{8}}$
$\quad+C\left\|\nabla\mathcal{G}_{\varphi_{1}}\Phi\right\|_{L^{2}(\Omega)}^{\frac{1}{4}}\left\|\nabla\Phi\right\|_{L^{2}(\Omega)}^{\frac{3}{4}}.$ (3.49)
@@
Concerning $I_{3}$, we observe from (1.20a) and (2.11) that
@@
$=\int_{\Omega}\nabla\mu_{2}\cdot\left(b(\varphi_{2})-b(\varphi_{1})\right)\nabla\mathcal{G}_{\varphi_{1}}\Phi\,{\rm d}x+\int_{\Omega}\nabla\mu_{2}\cdot b(\varphi_{1})\nabla\mathcal{G}_{\varphi_{1}}\Phi\,{\rm d}x$
$\quad-\int_{\Omega}\nabla\mu_{2}\cdot b(\varphi_{2})\nabla\mathcal{G}_{\varphi_{2}}\Phi\,{\rm d}x$
@@
$\leq C\left\|\nabla\mu_{2}\right\|_{L^{2}(\Omega)}\left\|\Phi\right\|_{L^{4}(\Omega)}\left\|\nabla\mathcal{G}_{\varphi_{1}}\Phi\right\|_{L^{4}(\Omega)}$
$\leq C\left\|\nabla\mu_{2}\right\|_{L^{2}(\Omega)}\left\|\nabla\mathcal{G}_{\varphi_{1}}\Phi\right\|_{L^{2}(\Omega)}^{\frac{1}{4}}\left\|\nabla\Phi\right\|_{L^{2}(\Omega)}^{\frac{3}{4}}$
×$\quad\times\left(\|\varphi_{1}\|_{H^{2}(\Omega)}^{\frac{1}{2}}\left\|\nabla\mathcal{G}_{\varphi_{1}}\Phi\right\|_{L^{2}(\Omega)}+\left\|\nabla\mathcal{G}_{\varphi_{1}}\Phi\right\|_{L^{2}(\Omega)}^{\frac{3}{4}}\left\|\nabla\Phi\right\|_{L^{2}(\Omega)}^{\frac{1}{4}}\right)$
$\leq C\left\|\nabla\mu_{2}\right\|_{L^{2}(\Omega)}\|\varphi_{1}\|_{H^{2}(\Omega)}^{\frac{1}{2}}\left\|\nabla\mathcal{G}_{\varphi_{1}}\Phi\right\|_{L^{2}(\Omega)}^{\frac{5}{4}}\left\|\nabla\Phi\right\|_{L^{2}(\Omega)}^{\frac{3}{4}}$
$\quad+C\left\|\nabla\mu_{2}\right\|_{L^{2}(\Omega)}\left\|\nabla\mathcal{G}_{\varphi_{1}}\Phi\right\|_{L^{2}(\Omega)}\left\|\nabla\Phi\right\|_{L^{2}(\Omega)}$
$\leq\frac{1}{8}\left\|\nabla\Phi\right\|_{L^{2}(\Omega)}^{2}+C\left(\left\|\nabla\mu_{2}\right\|_{L^{2}(\Omega)}^{\frac{8}{5}}\|\varphi_{1}\|_{H^{2}(\Omega)}^{\frac{4}{5}}+\left\|\nabla\mu_{2}\right\|_{L^{2}(\Omega)}^{2}\right)\left\|\nabla\mathcal{G}_{\varphi_{1}}\Phi\right\|_{L^{2}(\Omega)}^{2}$
$\leq\frac{1}{8}\left\|\nabla\Phi\right\|_{L^{2}(\Omega)}^{2}+C\left(\left\|\nabla\mu_{2}\right\|_{L^{2}(\Omega)}^{2}+\|\varphi_{1}\|_{H^{2}(\Omega)}^{4}\right)\left\|\nabla\mathcal{G}_{\varphi_{1}}\Phi\right\|_{L^{2}(\Omega)}^{2}.$ (3.51)
@@
$\frac{{\rm d}}{{\rm d}t}\left\|\sqrt{b(\varphi_{1})}\nabla\mathcal{G}_{\varphi_{1}}\Phi\right\|_{L^{2}(\Omega)}^{2}+\left\|\nabla\Phi\right\|_{L^{2}(\Omega)}^{2}\leq h(t)\left\|\sqrt{b(\varphi_{1})}\nabla\mathcal{G}_{\varphi_{1}}\Phi\right\|_{L^{2}(\Omega)}^{2}\!$ (3.53)
@@
$\begin{split}&\varphi\in L^{\infty}(0,\infty;W^{2,p}(\Omega)),\quad\partial_{t}\varphi\in L_{\operatorname{uloc}}^{2}(0,\infty;H^{1}(\Omega)),\\ &\varphi\in L^{\infty}(\Omega\times(0,\infty))\quad\text{such that}\quad|\varphi(x,t)|<1\ \text{a.e. }(x,t)\in\Omega\times(0,\infty),\\ &\mu\in L^{\infty}(0,\infty;H^{1}(\Omega))\cap L_{\operatorname{uloc}}^{4}([0,\infty);H^{2}(\Omega)),\quad F^{\prime}(\varphi)\in L^{\infty}(0,\infty;L^{p}(\Omega)),\end{split}$ (4.1)
@@
First of all, owing to [5, Theorem 2.2] (cf. the three dimensional case), there exists a unique strong solution $\varphi$ to problem (1.1a)-(1.1b) and (1.2) on a maximal time interval $[0,T_{\star})$. In particular, the solution $\varphi$ is the limit of an approximating sequence $\{\varphi_{\epsilon}\}$, where $\varphi_{\epsilon}$ is the solution to a suitable approximation problem associated with a regularized potential $F_{\epsilon}$ suitably converging to $F$ as $\epsilon\to 0$. Our aim is to show that $T_{\star}=+\infty$. More precisely, we will show that $\varphi$ satisfies (4.1). To this end, we will perform global-in-time estimates leading to (4.2)-(4.4) by directly working on the solution $\varphi$. However, a rigorous proof should be carried out by making the computations below at the approximation level $\varphi_{\epsilon}$, where also the initial datum is properly regularized (see the proof of [28, Theorem 4.1]), and by passing ultimately to the limit. Since this procedure is nowadays well established in the literature, we will leave this step to the interested reader.
@@
almost everywhere on $(0,T_{\star})$. The way to handle the first term on the right-hand side in (4.8) is well known (see (4.14) below). Instead, we need to find a suitable control for the second term in the right-hand side. In order to do so we present two different approaches.
@@
$\left\|\mathcal{G}_{\varphi}\partial_{t}\varphi\right\|_{H^{2}(\Omega)}\leq\left(\frac{2C}{\tau(1-\tau)}\right)^{\frac{1}{2(1-\tau)}}\left\|\nabla\varphi\right\|_{L^{2}(\Omega)}^{\frac{\tau}{1-\tau}}\left\|\varphi\right\|_{H^{2}(\Omega)}\left\|\nabla\mathcal{G}_{\varphi}\partial_{t}\varphi\right\|_{L^{2}(\Omega)}+\frac{C}{1-\tau}\left\|\partial_{t}\varphi\right\|_{L^{2}(\Omega)}\!.$
@@
$\begin{split}\left\|\nabla\mu\right\|^{4}_{L^{2(1+\tau)}(\Omega)}&\leq C(1+\tau)^{2}\|\nabla\mu\|_{L^{2}(\Omega)}^{\frac{4}{1+\tau}}\left\|\mathcal{G}_{\varphi}\partial_{t}\varphi\right\|_{H^{2}(\Omega)}^{\frac{4\tau}{1+\tau}}\\ &\leq C(1+\tau)^{2}\left(\frac{1}{\tau}\right)^{\frac{5\tau}{2(1+\tau)}}\|\nabla\varphi\|_{L^{2}(\Omega)}^{\frac{4\tau^{2}}{1-\tau^{2}}}\|\varphi\|_{H^{2}(\Omega)}^{\frac{4\tau}{1+\tau}}\|\nabla\mu\|_{L^{2}(\Omega)}^{4}\\ &\quad+C(1+\tau)^{2}\|\nabla\mu\|_{L^{2}(\Omega)}^{\frac{4}{1+\tau}}\|\partial_{t}\varphi\|_{L^{2}(\Omega)}^{\frac{4\tau}{1+\tau}}\\ &\leq C(1+\tau)^{2}\left(\frac{1}{\tau}\right)^{\frac{5\tau}{2(1+\tau)}}\|\nabla\varphi\|_{L^{2}(\Omega)}^{\frac{4\tau^{2}}{1-\tau^{2}}}\|\varphi\|_{H^{2}(\Omega)}^{\frac{4\tau}{1+\tau}}\|\nabla\mu\|_{L^{2}(\Omega)}^{4}\\ &\quad+\frac{2\tau}{1+\tau}\|\partial_{t}\varphi\|_{L^{2}(\Omega)}^{2}+\left(\frac{1-\tau}{1+\tau}\right)\left[C(1+\tau)^{2}\right]^{\frac{1+\tau}{1-\tau}}\ \|\nabla\mu\|_{L^{2}(\Omega)}^{\frac{4}{1-\tau}}.\end{split}$